The physician orders Medication B 80 mg IM now.

The vial of Medication B contains 50 mg /mL.

How many mL will the nurse prepare to administer?
Answer #1

ANSWER

* Amount of medication prepared for the patient.

       Formula.

         Amount(mL) of medication need to

         be administered = Amount desired(D) / Amount on hand(H) * Quantity(Q).

                  Amount desired (D) = 80 mg.

                  Amount on hand(H) = 50 mg.

                          Quantity (Q) = 1 mL.

         As per formula

                   Medication prepared by the

                    nurse to administer = D / H * Q.

                                                              = 80 / 50 * 1

                                                              = 1.6 * 1

                                                              = 1.6 mL.